A hydrocarbon Z with molecular mass 78 on combustion gave 3.385g of CO2 and 0.692g of H2O. Determine the molecular formula of Z [H=1, C=12, O=16]
1
Expert's answer
2019-07-01T03:55:47-0400

Z + O2 = CO2 + H2O


"moles \space of \space CO_2 = \frac {3.385 \space g}{44 \space g/mol} = 0.0769 mol"

"mass \space of \space carbon \space in \space Z= 12 g/mol \cdot 0.0769 mol= 0.923 g"


"moles \space of H_2O=\frac {0.692 g}{18g/mol}=0.0384 mol"

"mass \space of \space hydrogen \space in \space Z= 2 g/mol \cdot 0.0384 mol = 0.077 g"



Empirical formula of Z is (CH)n = 78 g/mol

13n =78

n=78/13

n= 6

Molecular formula of Z is C6H6